# Deep Submillimeter and Radio Observations in the SSA22 Field. III. Multiwavelength Identifications and Properties of the 850 -selected Submillimeter Galaxies

### v1

@@ III.1.1 Identification Result
III.2 + Identifications
Following Michałowski et al. (2012), we also employ IRAC 8 and red opt-NIR color as a method to search for potential counterparts. However, for reliability considerations, we only include those counterparts with 0.05 in the source catalog.
Similar to the MIR-FIR correlation mentioned earlier, exploration in the IRAC 8 band is often conducted to enhance the completeness of identifying submillimeter sources (Ashby et al., 2006; Michałowski et al., 2012; Koprowski, 2015; Koprowski et al., 2017; Michałowski et al., 2017; Zavala et al., 2018). SED analysis of SMGs suggested that their 8 emission primarily traces ongoing star formation (Ashby et al., 2006). The evolved stellar populations have been discovered in massive SMGs, the expected spectral signatures of evolved stellar populations should peak at rest-frame 2-3 (Koprowski, 2015), corresponding to observed 8 .
Some studies have used a NIR color to select high-redshift, extremely red galaxies, such as HIEROs (Wang et al., 2012), KIEROs (Wang et al., 2016), OIRTC (Chen et al., 2016). Similarly, due to the highly dust-obscured nature of SMGs often displaying red optical/NIR colors, many researchers have used color-selected galaxies to identify SMGs (Michałowski et al., 2012; Cowie et al., 2017; An et al., 2018; Zhang et al., 2022). Following Michałowski et al. (2012), we employed a red optical/NIR color criteria ( 2 & 23 mag) to identify potential SMGs counterparts. The detailed measurement description could be seen in the Section II.3.
SMGs typically exhibit redshifts in the range –3. In this redshift interval, the observed - and -bands correspond approximately to the rest-frame near-UV and -band, respectively, straddling the Balmer/4000Å break—a key spectral feature for galaxy characterization. The Balmer break arises at the Balmer limit (3645 ) due to bound-free absorption by hydrogen atoms with electrons in the energy level. This feature is weakened in hot OB stars because most hydrogen is ionized, but it is strongest in cooler A-type stars (– M⊙), whose effective temperatures (10,000 K) maximize the population of electrons in the state. This feature is particularly prominent in star-forming galaxies (Dudzevičiūtė et al., 2020b; Wilkins et al., 2023). Moreover, it becomes especially strong in galaxies with stellar populations of age 0.3 Gyr, when a large number of UV-bright stars begin to evolve off the main sequence (Bruzual and Charlot, 2003; Vikaeus et al., 2024). The 4000 Å break primarily driven by metal absorption lines from cool, evolved stars (Mo et al., 2010; Dudzevičiūtė et al., 2020b). Together, the Balmer/4000Å break serves as a powerful diagnostic of recent star formation history and is sensitive to dust attenuation, metallicity, and the Lyman continuum (LyC) escape fraction (Wilkins et al., 2023).
These selected galaxies are referred to as Extremely Red Galaxies (ERGs), which may stem from evolved stellar components or galaxies that have reddened significantly due to dust obscuration in star-forming galaxies or starbursts (or AGN) (Caputi et al., 2006; Schael, 2009; Mo et al., 2010). ERGs are believed to align on an evolutionary sequence with high-redshift SMGs and local massive ellipticals (Schael, 2009). In the redshift range of 1.5-3, 40-60% of high-mass ERGs are ULIRGs (Caputi et al., 2006). SMGs, characterized by severe dust obscuration and intense star formation, are expected to exhibit distinct optical/NIR features. Therefore, such selections seem to be a good criterion.
III.2.1 Identification Result and Reliablity
@@
Following the methodology of Michałowski et al. (2012), we cross-matched counterparts identified through the / method with those confirmed by Radio/ to evaluate the reliability of the / selection. In the deepest regions of the SCUBA-2 map with optimal multiband coverage, approximately 85% of / selected counterparts were independently confirmed by Radio/ methods (Figure 9), consistent with the 84% agreement reported by Schael (2009); Michałowski et al. (2012). This high confirmation rate demonstrates the robustness of the / method and reflects the physical correlations between different star formation tracers. We observed a declining match rate (to 70%) in the shallower outer regions of the SCUBA-2 map, which we attribute to decreasing multi-band detection sensitivity at the survey edges. This trend highlights the importance of uniform depth in multiwavelength ancillary data for reliable counterpart identification.
The high surface density of 8 and optical/NIR-red selected sources leads to 0.05 at separations of 2.4-2.8 (corresponding to 20-30 physical kpc at 2-3). Meanwhile, the observed positional offsets between / counterparts and SCUBA-2 centroids peak at 1.5 and 2.5. Apart from sources with line-of-sight associations, such distances make it less likely to misidentify physically unrelated sources near the centers of SCUBA-2 sources, provided that genuine SMGs exist at the SCUBA-2 centroid positions.
III.3 Multiplicity
Due to the SCUBA-2 850 beam size of approximately 14 (Zeng et al., 2024), interferometric observations indicate that the blending of SMGs in a single-dish source cannot be ignored, with this blending ratio being 20% (An et al., 2019). Furthermore, multiwavelength studies have shown that the likelihood of SCUBA-2 sources corresponding to multiple SMGs is approximately 20-40% (Chen et al., 2016; Hyun et al., 2023), and the ratio increase with the rising flux of submillimeter sources (Chen et al., 2016). These facts serve as a caveat to treat galaxies behind SCUBA-2 sources cautiously.
Overall, the multiplicity rate of SCUBA-2 sources is 26% (Figure 10). For SCUBA-2 sources with fluxes below 8 mJy, the multiplicity fraction ranges from approximately 10% to 40%. This fraction exhibits an increasing trend with higher flux densities, rising to 40-60% for sources exceeding 8 mJy. However, the limited number of bright sources results in significant uncertainties. When considering all sources above 8 mJy, the multiplicity fraction is 46%, consistent with the results reported by Chen et al. (2016).
@@ IV SED-fitting
IV.1 EAZY Photometeric Redshift
@@ IV.2 CIGALE SED Fitting
We perform SED fitting using Investigating GALaxy Emission (CIGALE version 2022.1; Noll et al., 2009; Boquien et al., 2019). By leveraging multi-wavelength photometry—from the optical/NIR through to radio—we derive key physical properties of our SMG sample. For galaxies lacking reliable or (Section IV.1), we additionally obtain thier far-infrared (FIR) photometric redshifts. This subsection describes the fitting configuration, and the resulting physical parameters.
@@
| attenuation powerlaw slope in the ISM | slope_ISM | -0.7 | |
| attenuation powerlaw slope in the BC | slope_BC | -1.3 | |
@@
| radiation field powerlaw slope | alpha | 2.0, 2.5, 3.0 | |
@@
| powerlaw slope of emission related to stellar | alpha_sf | 0.8 | |
| powerlaw slope of AGN radio emission | alpha_agn | 0.7 | |
| redshifting | reshift grid | redshift | 0.0001, |
@@ IV.2.1 Configuration
We adopt the sfh2exp model with a double exponential star formation history (SFH). By setting the mass fraction of the second burst, we determine whether the model experiences one or two star formation events, with the latter including one long-term star formation and one recent burst. Previous studies suggest that SMGs may undergo multiple intense star formation episodes (Michałowski et al., 2010). Both simulations and main-sequence models indicate that double SFHs provide a more likely description, regardless of whether galaxies evolve through mergers or in isolation (Michałowski et al., 2014; Speagle et al., 2014). An exponentially declining SFH also consistent with true values (Michałowski et al., 2014).
@@
The radio emission in most galaxies is predominantly powered by star formation activity, consisting of the non-thermal synchrotron radiation from relativistic electrons accelerated in supernova remnants, and thermal bremsstrahlung emission (free-free emission) from HII regions surrounding massive stars. While bremsstrahlung exhibits a flat spectrum, its contribution becomes negligible compared to synchrotron emission at frequencies below 30 GHz (Helou et al., 1985; Condon, 1992). The well-established radio-FIR correlation arises because both wavebands trace star formation activity (Section III.1). However, we allow for variations in the correlation coefficient to account for potential redshift evolution of the radio-FIR relation (Speagle et al., 2014), and possible radio excess caused by AGN contamination.
@@ V.1 Redshift Distribution
Figure 14 panels (a) and (d) show the redshift histogram and spatial distribution of the full SMG sample, respectively. The horizontal bars in panel (a) represent the comoving cosmological volume probed within each redshift bin. The data reveal a pronounced evolution in SMG number density: at z 4, the density increases sharply by approximately a factor of 6, while at 1-3 it rises more gradually, maintaining values between 1.78-3.16 10-5 cMpc-3. This two-phase evolutionary pattern suggests distinct regimes of SMG formation and growth across cosmic time (see Paper IV). Notably, this includes 115 newly identified SMGs (corresponding to 95 SCUBA-2 sources with fluxes 2.8-3.5 mJy) that we report for the first time, which exhibit a slightly lower median redshift of 1.9.
@@
V.1.1 Oversensity
C. C. Steidel, K. L. Adelberger, M. Dickinson, M. Giavalisco, M. Pettini, and M. Kellogg (1998); 107 identified a large-scale structure at in the central region of the SSA22 deep field using low-mass Lyman- emitters (LAEs). Our identified SMGs also exhibit significant overdensity within this redshift slice (Figure 14 (c)).
When considering the broader redshift interval – the overdense region contains 17 SMGs, corresponding to a 8.4 (7.8) overdensity for bin widths of 0.05 (0.1). The ADF22 survey (umehata2017; Umehata et al., 2018) contributes five additional counterparts, two of which lie within the overdense region (green shaded area), confirming that our identification methods already robustly trace this large-scale structure at 7 significance.
@@ V.1.1 Oversensity
Moreover, studies have shown that SMGs are frequently associated with other populations of galaxies within large-scale structures (Tamura et al., 2009; Umehata et al., 2014; Wang et al., 2025) and co-evolve with these cosmic environments (Umehata et al., 2019). These results highlight the role of large-scale structures in driving galaxy evolution, while conversely underscoring the reliability and potential of massive dusty galaxies as robust tracers of cosmic structure formation—particularly for mapping mass distributions in high-redshift overdense regions.
@@
| Porperties | Total sample | New detections | ‘MAIN’ sample | ‘SUPP’ sample | Redshift 2.0 | Redshift 2.0 | Optically faint | Optically bright |
| (221 SMGs) | (115 SMGs) | (155 SMGs) | (66 SMGs) | (110 SMGs) | (111 SMGs) | (73 SMGs) | (148 SMGs) | |
| Redshift | 2.00 | 1.90 | 2.07 | 1.83 | 1.25 | 2.70 | 2.20 | 1.75 |
| log [] | 11.19 | 11.20 | 11.23 | 11.03 | 10.93 | 11.32 | 11.29 | 11.14 |
| log SFR [ yr-1] | 2.22 | 2.07 | 2.25 | 2.06 | 1.58 | 2.55 | 2.25 | 2.12 |
| log []) | 12.35 | 12.25 | 12.40 | 12.32 | 11.91 | 12.69 | 12.42 | 12.31 |
| log [] | 9.29 | 9.22 | 9.22 | 9.48 | 9.35 | 9.19 | 9.35 | 9.27 |
| log / [] | 0.21 | 0.20 | 0.21 | 0.21 | 0.29 | 0.12 | 0.34 | 0.17 |
| [mag] | 3.09 | 2.82 | 3.08 | 3.16 | 3.10 | 3.06 | 3.87 | 2.82 |
| log Agem [yr] aaThe mass-weighted ages of the galaxies, computed as (Noll et al., 2009), are derived from CIGALE. | 8.75 | 9.04 | 8.75 | 9.05 | 9.17 | 8.62 | 8.91 | 8.75 |
| log sSFR [yr-1] | -9.05 | -9.18 | -9.08 | -9.03 | -9.37 | -8.83 | -9.15 | -9.00 |
| MS bbIdentification method. Most sources with ALMA counterparts are also recovered by our identification method. If additional identifications exist, the next most reliable method—ranked as —is indicated in parentheses. Owing to the superior astrometric precision of ALMA, we adopt ALMA-derived coordinates for source positions whenever available. | -0.16 | -0.21 | -0.19 | -0.03 | -0.20 | -0.15 | -0.27 | -0.08 |

### v2

@@ III.2 Identification Result
III.3 Identification Accuracy and Caveats
In this section, we provide a preliminary assessment of the reliability of our identification methods and offer important caveats for users of our catalog.
To evaluate the performance of our identification techniques, we cross-match our identified sources with two ALMA surveys in the SSA22 field: the ADF22 survey (including the ADF22 A/B fields; Umehata et al., 2017, 2018) and the deeper ADF22+ survey (Huang et al., 2025), as well as the ALMA 2 mm continuum survey (Cooper et al., 2022).
Figure 9 shows the footprint of the ADF22 survey. Focusing on SCUBA-2 sources detected at 4, the identification accuracies of our four methods—, , , —are 15/20 ( 75%), 5/6 ( 83%), 9/9 ( 100%), and 2/3 ( 67%), respectively. When including marginal detections in the range 3.5–4, the accuracies decrease modestly to 17/27, 5/6, 10/12, and 2/4. The high fidelity for 4 sources reflects the fact that the ADF22 fields lie within the region of most complete multiwavelength coverage, underscoring the critical importance of deep, multi-band data for reliable source identification. Survey depth also plays a role: the deeper ADF22+ observations confirmed three low-significance ADF22 A detections, one of which (ADF22+ source A28) corresponds to a component of the blended SCUBA-2 source SSA22.0070. This source was robustly identified by both our and methods.
Comparison with the ALMA 2 mm survey yields (Figure 10) lower—but still reasonable—identification accuracies: 13/19 ( 68%), 2/5 ( 43%), 9/12 ( 75%), and 8/13 ( 62%). We emphasize again that these accuracy estimates depend sensitively on both multiwavelength coverage and ALMA survey depth.
We further compare sources identified via and methods with those detected in radio and 24 . The matching fractions are consistent between the two methods (Figure 12), reaching 75–85% in the central region of the field (where SCUBA-2 map RMS 1 mJy). Both matching rates decline with increasing SCUBA-2 map noise, yielding overall matching fractions of 71% () and 67% ()—a direct consequence of the deepest and most complete multiwavelength coverage in the field center. Because IRAC 8 sources and color-selected objects have relatively high surface densities, positional association probabilities drop below 0.05 only at angular separations of 2.4–2.8. The median offsets between our and identifications and the SCUBA-2 centroids are 1.5 and 1.8, respectively—corresponding to physical scales of 10–15 kpc at 2. This is comparable to the pointing uncertainty of JCMT (Geach et al., 2017). Therefore, if SMGs reside near the centroid of the SCUBA-2 source, these two identification methods are unlikely to introduce significant misidentifications at such angular separations, aside from chance alignments along the line of sight.
The the 80% correspondence between sources identified via the / methods and those detected in radio or is remarkably high—indeed, a pleasantly surprising result. This strong overlap suggests an underlying physical connection: even when some of the IRAC 8 sources or color-selected objects identifications do not correspond to genuine SMGs, they are still likely to be galaxies undergoing active star formation. Moreover, as shown in the earlier matching analysis, the identification accuracy of the 8 method reaches 80%, demonstrating its considerable potential—comparable to the more established and techniques.
For the method, we perform additional tests to quantify the contamination from random alignments. First, within the ADF22 survey area, we mask all regions within 7 of any ALMA-detected SMG or SCUBA-2 source. We then place random positions and check whether a source lies within 2.78 ( 0.05). This yields a lower limit on the contamination rate due to chance superpositions. In this small masked region, we generate 200 random positions per trial and repeat the experiment 50 times, obtaining a false-match probability of 4.02 1.32%. Next, we expand the test to a larger rectangular area (R.A. = –, Dec. = +000700–+002800), which covers field center with the most complete multiwavelength coverage. Here, we draw 2000 random positions per trial over 10 repetitions, resulting in a false-match probability of 4.84 0.40%. Overall, thanks to our strict matching criterion ( 0.05), the expected contamination from random alignments is low—below 5%. However, it should be emphasized that this estimate represents a lower limit. In reality, the true contamination could be higher, because SCUBA-2 sources are preferentially located in overdense environments. Such clustering enhances the likelihood of finding a red, massive galaxy near a submillimeter position even without a physical association.
Deep, well-characterized ALMA follow-up surveys—such as ALESS (cunha2015) and AS2UDS (Stach et al., 2019; Dudzevičiūtė et al., 2020b)—contain only a small minority of sources at low redshift ( 1), with the vast majority lying on or above the star-forming main sequence. In contrast, our sample includes a non-negligible fraction of low-redshift or quiescent systems, regardless of the identification method used (see Figure 11). We acknowledge, however, that some of these apparently quiescent galaxies may be artifacts of SED fitting: When far-infrared constraints are weak or absent, it leads to increased scatter in derived physical properties and may result in spurious classifications as “quenched” systems. We briefly discuss this effect in Section V.2. We also emphasize that our multiwavelength identification approach inevitably leads to the loss of a fraction of high-redshift SMGs. To illustrate this, we apply simple selection criteria to the AS2UDS catalog (Dudzevičiūtė et al., 2020b), isolating subsamples with radio detections, MIPS 24 counterparts, IRAC 8 counterparts, or those satisfying the color cut 2 and 23 mag. The median redshifts of these subsamples are 2.45, 2.31, 2.58, and 2.29, respectively—systematically lower than that of the full AS2UDS sample ( 2.61). This trend underscores a key limitation: multiwavelength identification methods inherently risk missing the high-redshift SMGs. This bias is particularly pronounced when relying on 24 or optical/NIR data. In the former case, the incompleteness is likely driven by sensitivity limits at high redshift, while in the latter, severe dust attenuation renders many high-redshift SMGs undetectable in optical/NIR bands.
We issue the following cautions: (1) Sources in regions with incomplete multiwavelength coverage—approximately where the SCUBA-2 map RMS 1–1.2 mJy—should be treated with care. (2) Identifications based solely on the criterion require particular scrutiny due to higher contamination risk. (3) Sources with redshifts 1 or classified as quiescent are likely contaminants and should be interpreted cautiously. Secondly, it is important to emphasize that the current identification methods likely miss a fraction of genuine SMG counterparts, particularly at high redshifts. Lastly, while our identification methods recover the majority of the SCUBA-2 flux, this raises a secondary concern: misidentified counterparts could lead to incorrect deblending and flux allocation. While given the relatively low multiplicity rate and the small number of -only identifications, we consider this risk to be modest.
In summary, cross-matching with deep ALMA data indicates that the majority of our identified sources belong to the true SMG population. Consequently, our sample is suitable for robust statistical analyses—a conclusion we will reiterate in subsequent sections.
III.4 Multiplicity
ALMA interferometric observations have revealed that single-dish submillimeter sources are often blends of multiple SMGs (Karim et al., 2013). In a follow-up ALMA study of 716 UDS SCUBA-2 sources, Stach et al. (2018) found an overall multiplicity fraction of approximately 11%. This fraction increases with submillimeter flux density: it rises to 26% for sources with S 5 mJy and further to 44% for those brighter than 9 mJy. Among these blended systems, about 30% consist of physically associated components (Stach et al., 2018). The measured multiplicity also depends on observational depth, as fainter SMGs are commonly found to be blended even in ALMA “blank” maps (simpson2014). When such faint blends are included, the multiplicity fraction may increase by an additional 10–20% (Hodge et al., 2013; Simpson et al., 2015).
In SSA22 field, the overall multiplicity rate is %. In the “deep region” (Figure 1), where multiwavelength data are most complete, fraction rises to 4%, consistent with typical multiplicity rates reported in other multiwavelength identification studies ( 20–40%; Chen et al., 2016; An et al., 2019). Specifically, for SCUBA-2 sources with flux densities 8 mJy, the multiplicity rate is 24 3%. In contrast, for sources 8 mJy, the multiplicity rate increases significantly, reaching 47 13%. Within individual high-flux bins, the multiplicity fractions range from 40% to 60%, though these estimates carry large uncertainties due to small number statistics.
@@ IV SED-fitting
We present the photometric redshifts derived from EAZY (Section IV.1), a description of our CIGALE configuration (Section IV.2.1), and the resulting SED fitting outputs (Section IV.2.2).
IV.1 EAZY Photometric Redshift
@@ IV.2 CIGALE SED Fitting
We perform SED fitting using Investigating GALaxy Emission (CIGALE version 2022.1; Noll et al., 2009; Boquien et al., 2019). By leveraging multi-wavelength photometry—from the optical/NIR through to radio—we derive key physical properties of our SMG sample. For galaxies lacking reliable or (Section IV.1), we additionally obtain their far-infrared (FIR) photometric redshifts. This subsection describes the fitting configuration, and the resulting physical parameters.
@@
| attenuation power-law slope in the ISM | slope_ISM | -0.7 | |
| attenuation power-law slope in the BC | slope_BC | -1.3 | |
@@
| radiation field power-law slope | alpha | 2.0, 2.5, 3.0 | |
@@
| power-law slope of emission related to stellar | alpha_sf | 0.8 | |
| power-law slope of AGN radio emission | alpha_agn | 0.7 | |
| redshifting | redshift grid | redshift | 0.0001, |
@@ IV.2.1 Configuration
We adopt the sfh2exp model with a double exponential star formation history (SFH). By setting the mass fraction of the second burst, we determine whether the model experiences one or two star formation events, with the latter including one long-term star formation and one recent burst. Previous studies suggest that SMGs may undergo multiple intense star formation episodes, and an exponentially declining SFH also consistent with true values (Michałowski et al., 2014). Both simulations and main-sequence models indicate that double SFHs provide a more likely description, regardless of whether galaxies evolve through mergers or in isolation (Speagle et al., 2014).
@@
The radio emission in most galaxies is predominantly powered by star formation activity, consisting of the non-thermal synchrotron radiation from relativistic electrons accelerated in supernova remnants, and thermal bremsstrahlung emission (free-free emission) from HII regions surrounding massive stars. While bremsstrahlung exhibits a flat spectrum, its contribution becomes negligible compared to synchrotron emission at frequencies below 30 GHz (Helou et al., 1985; Condon, 1992). The well-established radio-FIR correlation arises because both wavebands trace star formation activity (Section III.1.1). However, we allow for variations in the correlation coefficient to account for potential redshift evolution of the radio-FIR relation (Speagle et al., 2014), and possible radio excess caused by AGN contamination.
@@ V RESULTS
In this section, we present and discuss the physical properties of our identified sample, including redshift (Section V.1), SFR/stellar mass (Section V.2), infrared luminosity (Section V.3), dust mass (Section V.4), and dust attenuation (Section V.5).
liao2024; uematsu2024 compared the physical properties derived from the CIGALE and MAGPHYS. They found that differences between dust modeling assumptions in these codes lead to modest systematic offsets in estimated dust masses and temperatures, as well as in stellar masses. Our CIGALE configuration closely follows that of liao2024; therefore, we expect good overall agreement between our derived physical parameters and those from MAGPHYS—except for dust mass, which we explicitly discuss in Section V.4. In the following sections, we will directly compare our estimates with published values.
@@ V.1 Redshift Distribution
Figure 17 panels (a) and (d) show the redshift histogram and spatial distribution of the full SMG sample, respectively. The horizontal bars in panel (a) represent the comoving cosmological volume probed within each redshift bin. The data reveal a pronounced evolution in SMG number density: at z 4, the density increases sharply by approximately a factor of 6, while at 1-3 it rises more gradually, maintaining values between 1.78-3.16 10-5 cMpc-3. This two-phase evolutionary pattern suggests distinct regimes of SMG formation and growth across cosmic time (see Paper IV of this series). Notably, this includes 115 newly identified SMGs (corresponding to 95 SCUBA-2 sources with fluxes 2.8-3.5 mJy) that we report for the first time, which exhibit a slightly lower median redshift of 1.9.
@@
V.1.1 Overdensity
C. C. Steidel, K. L. Adelberger, M. Dickinson, M. Giavalisco, M. Pettini, and M. Kellogg (1998); 113 identified a large-scale structure at in the central region of the SSA22 deep field using low-mass LAEs. Our identified SMGs also exhibit significant overdensity within this redshift slice (Figure 17 (c)).
When considering the broader redshift interval – the overdense region contains 17 SMGs, corresponding to a 8.4 (7.8) overdensity for bin widths of 0.05 (0.1). The ADF22 survey (Umehata et al., 2017, 2018) contributes five additional counterparts, two of which lie within the overdense region (green shaded area), confirming that our identification methods already robustly trace this large-scale structure at 7 significance.
@@ V.1.1 Overdensity
As shown in the figure 17 (e), Lyman-alpha Blobs (LABs)—objects exhibiting extreme galactic activity (Geach et al., 2005, 2014)—are co-spatial with our SMG sample within the large-scale structure of SSA22. Moreover, studies have shown that SMGs are frequently associated with other populations of galaxies within large-scale structures (Tamura et al., 2009; Umehata et al., 2014; Wang et al., 2025) and co-evolve with these cosmic environments (Umehata et al., 2019). These results highlight the role of large-scale structures in driving galaxy evolution, while conversely underscoring the potential of massive dusty galaxies as tracers of cosmic structure formation—particularly for mapping mass distributions in high-redshift overdense regions.
@@
| Porperties | Total sample | New detections aaCounterparts of submillimeter sources detected in our deeper observations. | ‘MAIN’ bbIdentification method. Most sources with ALMA counterparts are also recovered by our identification method. If additional identifications exist, the next most reliable method—ranked as —is indicated in parentheses. Owing to the superior astrometric precision of ALMA, we adopt ALMA-derived coordinates for source positions whenever available. | ‘SUPP’ | Conservative Sample ccTo minimize contamination, we adopt a conservative sample selection, retaining only sources with redshift 1 that have or counterparts. | 2.0 | 2.0 | Opt/NIR faint ddThe optically faint subsample consists of identifications with detections in fewer than four opt/NIR filters, whereas the optically bright subsample includes sources detected in four or more opt/NIR filters. | Opt/NIR bright |
| 221 ids | 115 ids | 155 ids | 66 ids | 133 ids | 110 ids | 111 ids | 73 ids | 148 ids | |
| Redshift | 2.00 | 1.90 | 2.07 | 1.83 | 2.24 | 1.25 | 2.70 | 2.20 | 1.75 |
| log Mstar [] | 11.19 | 11.20 | 11.23 | 11.03 | 11.28 | 10.93 | 11.32 | 11.29 | 11.14 |
| log SFR [ yr-1] | 2.22 | 2.07 | 2.25 | 2.06 | 2.38 | 1.58 | 2.55 | 2.25 | 2.12 |
| log LIR []) | 12.35 | 12.25 | 12.40 | 12.32 | 12.51 | 11.91 | 12.69 | 12.42 | 12.31 |
| log Mdust [] | 9.29 | 9.22 | 9.22 | 9.48 | 9.22 | 9.35 | 9.19 | 9.35 | 9.27 |
| log Mstar/ [] | 0.21 | 0.20 | 0.21 | 0.21 | 0.20 | 0.29 | 0.12 | 0.34 | 0.17 |
| [mag] | 3.09 | 2.82 | 3.08 | 3.16 | 3.09 | 3.10 | 3.06 | 3.87 | 2.82 |
| log Agem [yr] eeThe mass-weighted ages of the galaxies, computed as (Noll et al., 2009), are derived from CIGALE. | 8.75 | 9.04 | 8.75 | 9.05 | 8.70 | 9.17 | 8.62 | 8.91 | 8.75 |
| log sSFR [yr-1] | -9.05 | -9.18 | -9.08 | -9.03 | -9.00 | -9.37 | -8.83 | -9.15 | -9.00 |
| MS ffThe offset of SMGs from the main sequence, defined as MS = (sSFR/sSFR), where the main sequence relation is adopted from Speagle et al. (2014). See also Paper IV. | -0.16 | -0.21 | -0.19 | -0.03 | -0.20 | -0.20 | -0.15 | -0.27 | -0.08 |